OpenStudy (anonymous):

What is the simplified form of X^2-25/X+5 a. x - 5, with the restriction x ≠ -5 b. x + 5, with the restriction x ≠ 5 c. x - 5, with the restriction x ≠ 5 d. x + 5, with the restriction x ≠ -5 How do I find restrictions?

OpenStudy (anonymous):

Memorize \[ a^2 - b^2 = (a + b) (a - b) \] \[(x^2-25) /( x - 5) \] \[= ( x+ 5) ( x - 5))/( x - 5) \] cancel common terms \[= ( x+ 5) \]